OpenStudy (anonymous):

help me with this problem appliying logarithms applications solving this: the population of a city is growing at a rate proportional to its population. the population 20 years ago was 100 000 and today is it 150 000. find the population 20 years from now.

OpenStudy (anonymous):

dp/dt= k p

OpenStudy (anonymous):

whats dp/dt=kp

OpenStudy (anonymous):

lol, forget it Just remember this \[P=P_0 e^{rt}\]

OpenStudy (lilg132):

225,000

OpenStudy (anonymous):

yes that it i tryed to do it usind this formula but got a decimal number

OpenStudy (akshay_budhkar):

i like your reply imran.. "lol forget it" lolz :D

OpenStudy (lilg132):

imranmeah you didnt explain it at all

OpenStudy (anonymous):

so do you get 225 000

OpenStudy (anonymous):

15=10 e^20r find r r= 3/ 2e^20

OpenStudy (anonymous):

r=0.0202733 Ignore last result 100 000 e^(40(0.0202733))=225000...

OpenStudy (anonymous):

following the formula I got 1.5=e^20r am I on the right step

OpenStudy (lilg132):

can i show you an easier way?

OpenStudy (anonymous):

yes but using logs application

OpenStudy (lilg132):

we already know how much the population grew by in 20 years it increased from 100,000 to 150,000 now to work out the percentage it increased by in 20 years you divide the total number of population now by what it was 20 years ago so you do this 150,000 / 100,000 = 1.5 1.5 * 100 = 150% it grew by in 20 years 150% / 20 = 7.5% every year (this is irrelevant to this question but just showing how to get percentage for each year if you wanted to know) now we need to find out how much the population will be after another 20 years so the population now is 150,000 150,000 * 1.5 = 225,000 we dont need to change the percentage as it is trying to work out for the same amount of time dont forget to click good answer and this is how i worked it out

OpenStudy (anonymous):

hi I need touse the formula imranmeah wrote above I got different result

OpenStudy (anonymous):

can you t fo

OpenStudy (anonymous):

use the formula